OpenStudy (annah.bee):

The sum of the measures of the interior angles of a polygon is 1620°. How many sides does the polygon have?

OpenStudy (anonymous):

well they r gonna b greater than 90° angles so cant b less than 5...

OpenStudy (annah.bee):

ok

OpenStudy (anonymous):

got any choices??

OpenStudy (annah.bee):

idk 6 or 7 lol

OpenStudy (anonymous):

6 would put them @ 270 each

OpenStudy (anonymous):

so not that

OpenStudy (annah.bee):

its gonna be like 20 or something isnt it

OpenStudy (aihberkhan):

Okay to find the Sum you will have to do \((n-2)180\) So, \(1620 = (n-2)180\) Now just keep on solving \(n-2 = 9\) \(n = 11\) sides

OpenStudy (mathstudent55):

Use the formula: S = (n - 2)180 where S = the sum of the measures of the angles of a polygon, and n = number of sides You know S. Plug in the number for S, and solve for n, the number of sides.

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @annah.bee

OpenStudy (annah.bee):

thanks for everyones help.. i will fan all of you nd medal youall